a. Assume that the reserve ratio in a country is 2%.
(i) Use a formula to calculate the simple money multiplier, show all steps involved in the calculation.


(ii) Suppose that in 2019 customers deposit $200 into their banks. Based on the simple money multiplier calculated in part (i), calculate the total amount which the money supply in the banking system will eventually increase to. Show all steps involved in the calculation.

(iii) Assume the amount of funds deposited by customers in 2020 stands at $400. and the reserve ratio increases to 8%. Will the money supply in the banking system increase, decrease or stay the same from 2019 levels? If it increases or decreases, calculate by how much it will change. If it stays the same, explain why. Show all steps involved in any calculations.
1
Expert's answer
2020-05-08T15:14:26-0400

(i)"simple money multiplier=\frac{1}{0.02}=50"

(ii)"200\times\frac{1}{0.02}=10 000"

(iii)"400\times\frac{1}{0.08}=5000"

5 000-10 000=-5 000

money supply decreases as the reserve ratio has increased